OpenStudy (anonymous):

There are 8 section of seats in an auditorium. Each section contains at least 150 seats but not more than 200 seats. Which of the following could be the number of seat in this auditorium.

OpenStudy (anonymous):

If there are 150 in each section, we would multiply 8 times 150. If there are 200 seats in each section, we would multiply 8 times 200. The answer must be between those two answers.

OpenStudy (anonymous):

my multIPLE choice a.200 c.1100 e.1,700 b.1000 d.1300

OpenStudy (anonymous):

Only one of those falls between 8*150 and 8*200... which one is it?

OpenStudy (anonymous):

i multiply the answer isnt there

OpenStudy (anonymous):

what did you get when you multiplied?

OpenStudy (anonymous):

1200 and 1600

OpenStudy (anonymous):

yes! which of your answers is between those two numbers?

OpenStudy (anonymous):

o thanks

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

The least integer of a set of consecutive integer is -25.if the sum of these integer is 26 how many integer are in this set

OpenStudy (anonymous):

interesting... so it look like {-25, -24, -23,... ,-2, -1, 0, 1, 2, 3, 4,...} each positive will cancel each negative before we start to get a positive sum immediately after which number?

OpenStudy (anonymous):

wat do u mean

OpenStudy (anonymous):

well, if we start at zero and work our way out in both directions we get -1+1=0 they cancel -2+2=0 they cancel -3+3=0 they cancel they will keep cancelling like that until we get out to -25 and ...?

OpenStudy (anonymous):

Oh, sorry, I don't just give answers... we have to work our way there. -25 will cancel with 25, then the very next number will start to get us a positive sum.

OpenStudy (anonymous):

The very next number will be 26... which is the sum that we are looking for.

OpenStudy (anonymous):

26

OpenStudy (anonymous):

So the set has a bunch of numbers in it... 25 negatives that cancel with 25 positives there is also a zero in the middle and the 26 on the end finishes the set how many total numbers are in that bunch of numbers?

OpenStudy (anonymous):

51

OpenStudy (anonymous):

almost... did you miss the zero or the 26?

OpenStudy (anonymous):

52

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

o thanks for helping me :)

OpenStudy (anonymous):

You're welcome... I have to go now though... keep up the good work! :)

OpenStudy (anonymous):

ok cya